# GeoBenchX: Benchmarking LLMs for Multistep Geospatial Tasks

### v1

@@ 7 Conclusions
We have created a benchmark dataset for multistep geo-spatial task completion, introduced an evaluation framework, and evaluated the leading mainstream commercial LLMs on these tasks. The task profile and the evaluation approach reflect the practical requirements of GIS professionals and the use of LLM in a GIS-assistant mode.
Experiments indicate that commercial LLMs with function calling can successfully solve geospatial tasks even with casual, ambiguous instructions. Performance varies by task complexity, spatial thinking requirements, and model helpfulness bias.
Sonnet 3.5 and GPT-4o showed the best overall results and a balance between identifying unsolvable tasks and solving solvable ones. Anthropic models performed well on easier tasks, but their preference for providing solutions rather than rejecting tasks, high token usage, and price may disadvantage them for geospatial tasks. o3-mini and GPT-4o-mini overall had lower success rates, especially in easy tasks, though they excel at identifying unsolvable tasks and are affordable.
Limitations
@@
We did not tune prompts individually for each LLM; we evaluate only solution algorithm not the final response (for rejects and final numbers); the benchmark set size is moderate; we didn’t evaluate the variance of the responses (only the binomial confidence bounds).
Appendix A Prompts for solver agent
The system prompt is in Figure 1, the rules prompt is in Figure 2.
@@ Appendix A Prompts for solver agent
Figure 1: System prompt for solver agent.
@@
Figure 2: Rules prompt for solver agent.
Appendix B GeoSpatial Tools List
loaddata
@@ Appendix B GeoSpatial Tools List
loadgeodata
@@
getrasterpath
@@
getrasterdescription
@@
analyzerasteroverlap
@@
getvaluesfromrasterwithgeometries
@@
mergedataframes
@@
getuniquevalues
@@
filtercategorical
@@
filternumerical
@@
createbuffer
@@
makechoroplethmap
@@
filterpointsbyrastervalues
@@
selectfeaturesbyspatialrelationship
@@
calculatelinelengths
@@
calculatecolumns
@@
scalecolumnbyvalue
@@
makeheatmap
@@
visualizegeographies
@@
getcentroids
@@
plotcontourlines
Creates contour lines from raster data.
  • rasterpath: Path to the raster file
  • outputgeodataframename: Name for contour GeoDataFrame
  • interval: Interval between contour lines
  • min/maxvalue: Range for contours
  • plotresult: Whether to display plot
  • title: Plot title
generatecontoursdisplay
@@
makebivariatemap
@@
rejecttask

### v2

@@ 7. Conclusions
We have created a benchmark dataset for multistep geospatial tasks, introduced an evaluation framework and evaluated the leading mainstream commercial LLMs on these tasks.
The analysis indicates that the commercially available LLMs with function calling can be used successfully to solve geospatial tasks even when tasks are stated casually, with ambiguity, missing information, and minimal initial instructions. The performance of the models varies by the complexity of geospatial tasks, the degree of spatial thinking required to solve them, and how strongly the model is skewed to provide any answer at all rather than reject an unsolvable task.
In terms of comparative performance (see Table 7 and Figure 2), the OpenAI’s o4-mini showed the best results overall. Anthropic’s Claude Sonnet 3.5 ranked second overall by consistently placing in the top three for accuracy across multiple categories.
The three other models with top-3 accuracy in 2-3 categories showed clear trade-off between the ability to solve tasks and to identify unsolvable ones: Claude Sonnet 4 performs more than three times better on solvable tasks than on unsolvable ones thus losing in accuracy in separate task groups; GPT-4o and Gemini 2.5 Pro are better at correctly rejecting tasks than at solving them and thus have better accuracy on task groups.
The remaining models, GPT-4.1 and Google’s Gemini 2.0 Flash, are more balanced in rejecting and solving tasks and have lower accuracy across task groups.
Based on the cost of running the models, the best options for geospatial agents might be OpenAI’s o4-mini due to its combination of moderate costs and high accuracy, and Google’s Gemini 2.0 Flash as the most cost-efficient model while still providing decent results. Anthropic models performed well on easy tasks and solvable tasks in general; however, their preference for trying to provide any solution rather than rejecting, high token usage, and price might be disadvantages when using them for geospatial tasks.
8. Limitations
@@
We did not tune prompts individually for each LLM; we did not provide human feedback to the model to guide solution; we evaluate only solution algorithm not the final response (for rejects and final numbers); the benchmark set size is moderate; we did not evaluate the variance of the responses (only the binomial confidence bounds).
Appendix A: Prompts for solver agent
System prompt is in Figure 3, rules prompt is in Figure 4.
@@
Figure 3. System prompt for solver agent.
@@
Figure 4. Rules prompt for solver agent.
Appendix B: GeoSpatial Tools List
loaddata:
@@
    state: Current graph state containing (Geo)DataFrames and messages
loadgeodata:
@@
    state: Current graph state containing (Geo)DataFrames and messages
getrasterpath:
@@
    state: Current graph state containing (Geo)DataFrames and messages
getrasterdescription:
@@
    state: Current graph state containing (Geo)DataFrames and messages
analyzerasteroverlap:
@@
    state: Current graph state containing (Geo)DataFrames and messages
getvaluesfromrasterwithgeometries:
@@
    state: Current graph state containing (Geo)DataFrames and messages
mergedataframes:
@@
    state: Current graph state containing (Geo)DataFrames and messages
getuniquevalues:
@@
    state: Current graph state containing (Geo)DataFrames and messages
filtercategorical:
@@
    state: Current graph state containing (Geo)DataFrames and messages
filternumerical:
@@
    state: Current graph state containing (Geo)DataFrames and messages
createbuffer:
@@
    state: Current graph state containing (Geo)DataFrames and messages
makechoroplethmap:
@@
    state: Current graph state containing (Geo)DataFrames and messages
filterpointsbyrastervalues:
@@
    state: Current graph state containing (Geo)DataFrames and messages
selectfeaturesbyspatialrelationship:
@@
    state: Current graph state containing (Geo)DataFrames and messages
calculatelinelengths:
@@
    state: Current graph state containing (Geo)DataFrames and messages
calculatecolumns:
@@
    state: Current graph state containing (Geo)DataFrames and messages
scalecolumnbyvalue:
@@
    state: Current graph state containing (Geo)DataFrames and messages
makeheatmap:
@@
    state: Current graph state containing (Geo)DataFrames and messages
visualizegeographies:
@@
    state: Current graph state containing (Geo)DataFrames and messages
getcentroids:
@@
    state: Current graph state containing (Geo)DataFrames and messages
generatecontoursdisplay:
@@
    plottitle: Title for the plot
    cleanupfiles: Whether to remove temporary files after processing
    colormap: Matplotlib colormap name
    figsize: Figure size as JSON array [width, height]
    addcolorbar: Whether to add a colorbar to the plot
    plotbackground: Whether to display raster as background
    state: Current graph state containing (Geo)DataFrames and messages
makebivariatemap:
@@
    state: Current graph state containing (Geo)DataFrames and messages
rejecttask:
@@
    state: Current graph state containing (Geo)DataFrames and messages